Laser Absorption Spectroscopy of TiCl^{+} : Analysis of the [17.9]^{3}Δ-X^{3}Φ(0,0) Band

\underline{L. A. Kaledin}, J. P. Parrish, and M. C. Heaven, Department of Chemistry, Emory University, Atlanta, GA 30322.

The laser absorption spectrum of the TiCl^{+} molecular ion was observed for the first time in a hollow cathode discharge. The discharge was driven at 25 kHz at powers in the range of 80-100 W. Peak currents of 0.5 A occured during the conducting half-cycles. Flowing samples of TiCl_{4} in Ar were discharged. 235 lines of the 48Ti^{35}Cl^{+} [17.9]^{3}Δ-X^{3}Φ (0,0) band and 122 lines of the ^{48}Ti^{37}Cl^{+} [17.9]^{3}Δ-X^{3}Φ (0,0) band were recorded at resolution 0.02 cm^{−1}. This band system was observed previously ^{a} in the emission spectrum of TiCl^{+} but the electronic assignments for the lower and upper states of the transition are revised here. Rotational constants (in cm^{−1}) and internuclear distances r_{0} (in nm) for the [17.9]^{3}Δ (v=0) and X^{3}Φ (v=0) states of ^{48}Ti^{35}Cl^{+} and 48Ti^{37}Cl^{+} are respectively:

\begin{matrix}B′_{0} = & 0.178692(16), & r′_{0} = & 0.2160(1), & B′′_{0} = & 0.174212(16), & r′′_{0} = & 0.2188(1), \\ B′_{0} = & 0.173076(18), & r′_{0} = & 0.2160(1), & B′′_{0} = & 0.168742(18), & r′′_{0} = & 0.2188(1). \end{matrix}

These results are in excellent agreement with the predictions based on the ligand field theory models ^{b}. The symmetry for the ground state of the TiCl^{+} molecular ion is determined to be X^{3}Φ. Work supported by AFOSR under grant F19628-90-C-025.

aW. J. Balfour and K. S. Chandrasekhar, J. Mol. Spectrosc. 139, 245-252 (1990).

bL. A. Kaledin, J. E. McCord, and M. C. Heaven, J. Mol. Spectrosc. 173, 499-509 (1995).